# Infinite-Time Mean Field FBSDEs and Viscosity Solutions to Elliptic Master Equations

### 27 Mar 2026 v3

@@ Abstract
This paper presents a further investigation of the properties of infinite-time mean field FBSDEs and elliptic master equations, which were introduced in [14] as mathematical tools for solving discounted infinite-time mean field games. By establishing the continuous dependence of the FBSDE solutions on their initial values, we prove the flow property of the mean field FBSDEs. And then, we prove that, at the Nash equilibrium, the value function of the representative player constitutes a viscosity solution to the corresponding elliptic master equation. Furthermore, we establish the uniqueness of classical solutions to the elliptic master equation under displacement monotonicity and certain growth assumptions.
@@ 1 Introduction
Forward-backward stochastic differential equations (FBSDEs) also serve as a powerful tool for the study of mean field games. The theory of general nonlinear BSDEs was pioneered by Pardoux and Peng [11, 12] in the early 1990s. [13] studied the infinite-time FBSDEs and established connections with quasilinear elliptic PDEs. Recently, [1] extended this framework to the McKean-Vlasov FBSDEs. This paper investigates the properties of infinite-time mean field FBSDEs, and employs them to address the existence of viscosity solutions for elliptic master equations.
@@
$\min_{\beta}J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right].$ (1.2)
@@
is a viscosity solution to elliptic master equation (1.9), and satisfies the relationship
$\partial_{x}V(x,\mu)=Y_{0}^{x,\xi}.$ (1.15)
Finally, we address the uniqueness of classical solutions to the elliptic master equation. We introduce a displacement monotonicity condition (see [5]), which states that for any square integrable random variables $\xi_{1},\xi_{2}$, the following inequality holds:
@@
Under suitable growth conditions and this displacement monotonicity condition, we establish the uniqueness of classical solutions to the elliptic master equation.
@@ 2 Preliminaries
$\mathbb{E}\left[\int_{t_{0}}^{\infty}e^{-Kt}|v_{t}|^{2}dt\right]<+\infty.$ (2.2)
To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$ and define the exponentially weighted $L^{2}$ norm
@@ 4.1 Differentiability of mean field FBSDEs
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(Y^{x+\delta,\xi}-Y^{x,\xi}\right)-\nabla Y^{x,\xi}\right\rVert_{r}=0.$
@@
and $H_{xx}^{\delta},H_{yy}^{\delta}$ are defined similarly. Finally, define
@@ 4.2 Viscosity solution to elliptic master equation
Remark 4.8
For the distribution-independent case, in which $b(x,\mu,a)$ and $f(x,\mu,a)$ reduce to $b(x,a)$ and $f(x,a)$, and consequently the Hamiltonian $H(x,\mu,y)$ becomes $H(x,y)$, we can obtain a stronger conclusion. Considering the unique $L_{r}^{2}$-solution to the reduced mean field FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{x}=\partial_{y}H(X_{t}^{x,\xi},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x,\end{cases}$ (4.59)
@@
where $f(x,\hat{\alpha}(x,y))$ can be identified with $H(x,y)-\partial_{y}H(x,y)\cdot y$. From the preceding derivation, it follows that $V(x)$ is a classical solution to the equation
@@ 4.3 Uniqueness of classical solutions to elliptic master equation
$U_{4}(x,\mu)=-x^{2}+3x\bar{\mu}-\frac{3}{2}(\bar{\mu})^{2}-\frac{1}{2}.$
This demonstrates that the classical solutions to the elliptic master equation are not unique in general. We will now prove that if we consider only the smooth solutions satisfying certain growth and monotonicity conditions, the elliptic master equation admits at most one regular solution.
@@
    $U$ is of at most quadratic growth. $\partial_{x}U$ is Lipschitz continuous, and its derivatives, $\partial_{xx}U$, $\partial_{xxx}U$, $\partial_{\mu}\partial_{x}U$ and $\partial_{\tilde{x}\mu}\partial_{x}U$ are of at most linear growth.
@@
for any $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$, which shows that the value function is displacement monotone.
@@
There exists at most one regular solution to the elliptic master equation (1.9).
@@
Next, we apply Itô’s formula to $M_{t}$ on $[0,T]$. Noticing that $\partial_{y}H(x,\mu,y)$ is Lipschitz continuous and considering the linear growth condition of the derivatives of $\partial_{x}\bar{U}$, we get that

### 27 Apr 2026 v4

# Infinite-Time Mean Field FBSDEs and the Associated Elliptic Master Equations

@@ Abstract
This paper presents a further investigation of the properties of infinite-time mean field forward-backward stochastic differential equations (FBSDEs) and the associated elliptic master equations, which were introduced in [18] as mathematical tools for solving discounted infinite-time mean field games. By establishing the continuous dependence of the FBSDE solutions on their initial values, we prove the flow property of the mean field FBSDEs. And then, we prove that, at the Nash equilibrium, the value function of the representative player constitutes a viscosity solution to the corresponding elliptic master equation. In particular, when the coefficients of the equations are distribution-independent, we construct a classical solution to the elliptic partial differential equation (PDE) via fully coupled infinite-time FBSDEs. Furthermore, for classical solutions possessing displacement monotonicity and certain growth conditions, we establish their uniqueness for the elliptic master equation.
@@ 1 Introduction
Forward-backward stochastic differential equations (FBSDEs) also serve as a powerful tool for the study of mean field games. The theory of general nonlinear backward stochastic differential equations (BSDEs) was pioneered by Pardoux and Peng [13, 14] in the early 1990s. It’s well known that BSDEs jointly with forward classical stochastic differential equations (SDEs) can provide a probabilistic interpretation of the solution of a class of second order semilinear parabolic and elliptic partial differential equations (PDEs), which generalizes the Feynman-Kac formula to nonlinear cases and plays an important role in the field of stochastic optimization. Here we refer the reader to [15, 12, 5]. For fully coupled infinite-time FBSDEs, [16] established the well-posedness, and [17] investigated the problem in a broader solution space and connected it to quasilinear elliptic PDEs. Subsequently, [1] extended this framework to McKean-Vlasov FBSDEs and used it to solve infinite-time mean field control problems and mean field games. This paper investigates the properties of infinite-time mean field FBSDEs, a class of fully coupled McKean-Vlasov infinite-time FBSDEs which were introduced in [18], and employs them to address the existence of viscosity solutions and the uniqueness of certain classical solutions for the associated elliptic master equations. More notably, for the distribution-independent setting, we construct a classical solution to the elliptic PDE using fully coupled infinite-time FBSDEs.
@@
$\min_{\beta}J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right].$ (1.2)
@@
satisfies the relationship
$\partial_{x}V(x,\mu)=Y_{0}^{x,\xi},$ (1.15)
and constitutes a viscosity solution to elliptic master equation (1.9). It must be emphasized that, due to the lack of a proof for the uniqueness of viscosity solutions, we can only assert that if the value function is sufficiently smooth, then it constitutes a classical solution to the elliptic master equation. Turning to the distribution-independent setting, however, the value function directly constitutes a classical solution to an elliptic PDE. Consider the solution to the reduced mean field FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{x}=\partial_{y}H(X_{t}^{x},Y_{t}^{x})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\left[\partial_{x}{H}(X_{t}^{x},Y_{t}^{x})-rY_{t}^{x}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x,\end{cases}$ (1.16)
and the value function
where $f(x,\hat{\alpha}(x,y))$ is the distribution-independent version of $f(x,\mu,\hat{\alpha}(x,y))$. We derive that $V(x)$ is a classical solution to the equation
$rU(x)=H(x,\partial_{x}U(x))+\frac{1}{2}\partial_{xx}U(x).$ (1.18)
The result can be naturally extended to multiple dimensions, which enables us to construct a solution to the nonlinear elliptic PDE
$rU(x)=H(x,\partial_{x}U(x))+\frac{1}{2}\Delta U(x),\quad x\in\mathbb{R}^{d}$ (1.19)
using fully coupled infinite-time FBSDEs. Here, $\Delta$ denotes the Laplacian operator. Finally, we address the uniqueness of classical solutions to the elliptic master equation. An example in [18] illustrates that classical solutions to the elliptic master equation are not unique; however, at most one of them can be used to construct a Nash equilibrium. We introduce the displacement monotonicity condition (see [6]) for the solutions, which states that for any square integrable random variables $\xi_{1},\xi_{2}$, the following inequality holds:
@@
When the setting is independent of the distribution, the condition reduces to $(x_{1}-x_{2})(\partial_{x}U(x_{1})-\partial_{x}(x_{2}))\geq 0$ for all $x_{1},x_{2}\in\mathbb{R}$, which means that $U$ is a convex function. For classical solutions with this displacement monotonicity and under suitable growth conditions, we establish their uniqueness for the elliptic master equation.
@@ 2 Preliminaries
$\mathbb{E}\left[\int_{t_{0}}^{\infty}\mathrm{e}^{-Kt}|v_{t}|^{2}dt\right]<+\infty.$ (2.2)
To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$ and define the exponentially weighted $L^{2}$ norm $\lVert\cdot\rVert_{K}$ by
@@ 4.1 Differentiability of mean field FBSDEs
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(Y^{x+\delta,\xi}-Y^{x,\xi}\right)-\nabla Y^{x,\xi}\right\rVert_{r}=0,$
$\lim_{\delta\to 0}\left\lvert\frac{1}{\delta}\left(Y^{x+\delta,\xi}_{0}-Y^{x,\xi}_{0}\right)-\nabla Y^{x,\xi}_{0}\right\rvert=0.$
@@
and $H_{xx}^{\delta},H_{yy}^{\delta}$ are defined similarly. Define
@@
Then, we have
$\lim_{\delta\to 0}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left((X_{t}^{\delta})^{2}+(Y_{t}^{\delta})\right)^{2}\mathrm{d}t\right]=0.$ (4.21)
At last, we apply Itô’s formula to $\mathrm{e}^{-rt}(Y^{\delta}_{t})^{2}$. It’s easy to derive that
$\lim_{\delta\to 0}\left\lvert Y_{0}^{\delta}\right\rvert^{2}=0.$ (4.22)
@@ 4.2 Viscosity solution to elliptic master equation
In Theorem 4.2 and 4.4, we prove the existence of both $\partial_{x}V(x,\mu)$ and $\partial_{xx}V(x,\mu)$ for the value function $V(x,\mu)$. This indicates that the proof of the viscosity solution deals exclusively with the distributional component. For the distribution-independent case, in which $b(x,\mu,a)$ and $f(x,\mu,a)$ reduce to $b(x,a)$ and $f(x,a)$, and consequently the Hamiltonian $H(x,\mu,y)$ becomes $H(x,y)$, we can obtain a stronger conclusion. Considering the unique $L_{r}^{2}$-solution to the reduced mean field FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{x}=\partial_{y}H(X_{t}^{x},Y_{t}^{x})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\left[\partial_{x}{H}(X_{t}^{x},Y_{t}^{x})-rY_{t}^{x}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x,\end{cases}$ (4.61)
@@
where $f(x,\hat{\alpha}(x,y))$ can be identified with $H(x,y)-\partial_{y}H(x,y)\cdot y$. From the preceding derivation, we have the following proposition.
Proposition 4.8
Suppose that the reduced Hamiltonian $H(x,y)$ satisfies Assumption 3.1 and 4.1. We have that $V(x)$ defined in (4.62) is a classical solution to the equation
@@ 4.3 Uniqueness of classical solutions to elliptic master equation
$U_{4}(x,\mu)=-x^{2}+3x\bar{\mu}-\frac{3}{2}(\bar{\mu})^{2}-\frac{1}{2},$
where we denote by $\bar{\mu}$ the expectation of the measure $\mu$. This demonstrates that the classical solutions to the elliptic master equation are not unique in general. We will now prove that if we consider only the smooth solutions satisfying certain growth and monotonicity conditions, the elliptic master equation admits at most one regular solution.
@@
    $U$ is of at most quadratic growth. $\partial_{x}U$ is Lipschitz continuous, and its derivatives, $\partial_{xx}U$ and $\partial_{\mu}\partial_{x}U$ are bounded, $\partial_{xxx}U$ and $\partial_{\tilde{x}\mu}\partial_{x}U$ are of at most linear growth.
@@
for any $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$, which shows that the value function is displacement monotone. We particularly remark that for the distribution-independent case, as shown in Proposition 4.8, the displacement monotonicity condition reduces to the convexity condition.
@@
Under Assumption 3.1, there exists at most one regular solution to the elliptic master equation (1.9).
@@
Next, we apply Itô’s formula to $M_{t}$ on $[0,T]$ and take expectation. We get that
$\mathbb{E}[M_{T}]=\mathbb{E}\Bigg[$ $\int_{0}^{T}\mathrm{e}^{-rt}\Bigg(-r(\bar{X}^{\xi}_{t}-X^{\xi}_{t})(Y^{\xi}_{t}-\partial_{x}\bar{U}(X^{\xi}_{t},{\cal L}_{X^{\xi}_{t}}))$ (4.75)
$+\bigl(\partial_{y}H(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}},\bar{Y}^{\xi}_{t})-\partial_{y}H(X_{t}^{\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{\xi})\bigr)(Y^{\xi}_{t}-\partial_{x}\bar{U}(X^{\xi}_{t},{\cal L}_{X^{\xi}_{t}}))$
$+(\bar{X}^{\xi}_{t}-X^{\xi}_{t})\bigl(-\partial_{x}H(\bar{X}^{\xi}_{t},\mathcal{L}_{\bar{X}^{\xi}_{t}},\bar{Y}^{\xi}_{t})+r\bar{Y}^{\xi}_{t}\bigr)$
$-(\bar{X}^{\xi}_{t}-X^{\xi}_{t})\bigg(\partial_{xx}\bar{U}(X^{\xi}_{t},{\cal L}_{X^{\xi}_{t}})\partial_{y}H(X_{t}^{\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{\xi})+\frac{1}{2}\partial_{xxx}\bar{U}(X^{\xi}_{t},{\cal L}_{X^{\xi}_{t}})$
$\qquad+\tilde{\mathbb{E}}_{{\cal F}_{t}}\Big[\partial_{\mu}\partial_{x}\bar{U}(X^{\xi}_{t},{\cal L}_{X^{\xi}_{t}},\tilde{X}^{\xi}_{t})\partial_{y}H(\tilde{X}_{t}^{\xi},{\cal L}_{X_{t}^{\xi}},\tilde{Y}_{t}^{\xi})$
$\qquad\qquad+\frac{1}{2}\partial_{\tilde{x}\mu}\partial_{x}\bar{U}(X^{\xi}_{t},{\cal L}_{X^{\xi}_{t}},\tilde{X}^{\xi}_{t})\Big]\bigg)\Bigg)\mathrm{d}t\Bigg].$
Noticing that $\partial_{y}H(x,\mu,y)$ is Lipschitz continuous and considering the growth condition of $\partial_{x}\bar{U}$, we get that